OpenStudy (goformit100):

In a cricket match, five batsmen A, B, C, D and E scored an average of 36 runs. D Scored 5 more than E; E scored 8 fewer than A; B scored as many as D and E combined; and B and C scored 107 between them. How many runs did E score ? A. 62 B. 45 C. 28 D. 20

OpenStudy (goformit100):

@Noemi95

OpenStudy (anonymous):

so Aruns + Bruns + Cruns+Druns +Eruns/5=36. Alot of subbing here E+5=D A-8=E B=D+E B+C=107. It looks like the common variable we can get here is E. So sub till you get the original equation all in E.

OpenStudy (goformit100):

How ?

OpenStudy (anonymous):

ok. (A) E+8 + (b+c) 107 + (D) E+5 +E/5=36

OpenStudy (anonymous):

bueno?